# LLM-based Relevance Assessment for Web-Scale Search Evaluation at Pinterest

### v1

@@ 1. Introduction
Relevance evaluation typically relies on human annotations. However, this approach suffers from high costs, long turnaround time, and limited scalability, which can ultimately compromise the quality of relevance metrics. In recent years, with the advancement of Large Language Models (LLMs), there is a growing interest in automating relevance assessment in the Information Retrieval (IR) community. There have been several works exploring the feasibility of using LLMs to supplement laborious human labeling efforts (Faggioli et al., 2023; Thomas et al., 2024; Upadhyay et al., 2024b).
@@ 2.1. Search Relevance Modeling
For LLM-based search relevance modeling, there are two types of architectures commonly used in practice: bi-encoders and cross-encoders. Bi-encoder models encode queries and documents separately into a common dense space and then score their relevance using vector dot-product or cosine similarity. These models are typically trained by minimizing the contrastive loss with in-batch negative sampling (Karpukhin et al., 2020; Ma et al., 2023). While bi-encoder models are efficient for retrieval tasks, they are limited in their ability to capture complex interactions between queries and documents. In contrast, cross-encoder models encode queries and documents jointly, hence can better capture the interaction between queries and documents. Various loss functions are used in the literature for training cross-encoder relevance models, such as pointwise classification loss (Nogueira et al., 2020), pairwise or listwise ranking loss (Zhuang et al., 2023), and contrastive loss (Ma et al., 2023). In this work, we adopt a cross-encoder architecture and fine-tune the LLMs by minimizing the point-wise multi-class classification loss. The fine-tuned LLMs are then applied to whole-page relevance assessment at Pinterest Search.
@@ 2.2. LLMs for Relevance Judgments
Recent studies have shown the potential of using LLMs for automating relevance assessments in Information Retrieval (IR), as part of the broader “LLM-as-a-judge” paradigm (Gu et al., 2024; Zheng et al., 2023). Initial explorations by Faggioli et al. (2023) examined the application of LLMs to relevance judgment, providing insights into human-LLM collaboration. Subsequent work by Thomas et al. (2024) and Upadhyay et al. (2024b) further demonstrated the capability of LLMs to accurately predict searchers’ preferences and generate human-quality relevance labels. Most current research investigates various prompting techniques for LLM-based relevance labeling, including zero-shot (Upadhyay et al., 2024b; Arabzadeh and Clarke, 2025; Upadhyay et al., 2024a) and few-shot (MacAvaney and Soldaini, 2023; Pires et al., 2025) approaches, with a primary focus on evaluating their alignment with human labeling. Later, Meng et al. (2025) and Abbasiantaeb et al. (2024) showed that fine-tuning open-source LLMs on human-labeled relevance judgments is crucial for obtaining more reliable relevance predictions, suggesting that fine-tuning much smaller LLMs can yield more effective relevance prediction than few-shot prompting with significantly larger models.
Unlike most existing work that primarily compares different LLMs and prompting techniques (Meng et al., 2025; Abbasiantaeb et al., 2024) or explores human-LLM collaboration (Shankar et al., 2024), our work presents a real-world application of LLM-based relevance assessment to measure relevance change in online experiments at Pinterest Search. We discuss the practical considerations and validation steps, and highlight the opportunities for improving relevance measurement driven by this approach.
@@ 3.2. Fine-tuned LLMs as Relevance Model
Figure 1. The cross-encoder architecture in the relevance teacher model. Take the encoder language models (e.g., BERT-based models) for illustration.
[…]
The figure illustrates the cross-encoder architecture in the relevance teacher model. We take the encoder language models (e.g., BERT-based models) for example.
Figure 2. Components of LLM-based relevance measurement at Pinterest Search.
[…]
Components of LLM-based relevance measurement at Pinterest Search: data collection, relevance labeling, and metrics calculation.
To effectively represent each Pin for relevance prediction, we leverage a comprehensive set of textual features, including Pin titles and descriptions, BLIP image captions (Li et al., 2022), linked page titles and descriptions, user-curated board titles where the Pin has been saved, and highly-engaged queries associated with the Pin. These features together form a robust text representation crucial for accurate relevance assessment.
We experiment with various pre-trained language models, including multilingual BERTbase (Devlin et al., 2018), T5base (Raffel et al., 2020), mDeBERTaV3base (He et al., 2021), XLM-RoBERTalarge (Conneau et al., 2019), and Llama-3-8B (AI@Meta, 2024). The comparative performance for these models and ablation studies on Pin text features can be found in Wang et al. (2024). We then use the fine-tuned model to generate 5-dimensional relevance scores and use the label corresponding to the highest score (argmax) for relevance assessment.
@@ 3.3. Stratified Sampling Design
Stratification plays an important role in sampling-based measurement. First, stratification ensures the sample population is representative of the whole population. In addition, if the strata are chosen such that each stratum is relatively homogeneous, variance reduction can be achieved (Miratrix et al., 2013). From a statistical perspective, consider the whole-page relevance score as a random variable $Y$ with mean $\mu$ and variance $\sigma^{2}$. The variance of the sample mean $\bar{Y}$ is given by $V(\bar{Y})=\sigma^{2}/N$, where $N$ is the sample size. If $K$ strata are chosen such that each stratum has a different mean, $V(\bar{Y})$ can be decomposed into within-strata and between-strata variance (see Equation 1), the latter of which can be removed due to stratified sampling (Xie and Aurisset, 2016),
@@
To determine the query strata, we evaluated multiple strata choices. While BERTopic (Grootendorst, 2022) produced the largest variance reduction, we adopted the in-house query-to-interest model based on DistilBERT (Sanh et al., 2019), due to ease of integration with existing systems and scalability.
@@ 3.4. Relevance Measurement with LLMs
To measure the relevance impact of an A/B experiment on search ranking, we take a stratified sample of paired search queries from control and treatment experiment groups, ensuring that the sample is representative of overall user usage (Thompson, 2012). The use of paired samples blocks between-query differences, an important source of variation in experiment measurement.
@@
(2) $sDCG@K=\frac{\sum_{k=1}^{K}L/log_{2}(1+k)}{\sum_{k=1}^{K}5/log_{2}(1+k)},\text{ }L\in\{1,2,3,4,5\}.$
Lastly, we calculate heterogeneous effects by query popularity and query interest (e.g. beauty, women’s fashion, art, etc), utilizing a Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995) to control the false discovery rate. The LLM-based relevance measurement procedure at Pinterest Search is illustrated in Figure 2.
@@ 4.1. RQ1: Alignment with Human Labels
Following the established best practices in literature (Meng et al., 2025; Abbasiantaeb et al., 2024) to measure alignment between LLMs and human labels, we compute and report the rank-based correlation Kendall’s $\tau$ and Spearman’s $\rho$ to assess the correlation between the two rankings at query-level sDCG@K metric. To understand the performance on queries with different popularity, we also categorize the queries into 4 popularity segments based on search volume (Jain and Mishne, 2010): head, torso, tail, and single [^2] Single queries are defined as those that receive fewer than 10 searches.. The results are summarized in Table 1. For all query popularity segments, we achieve Kendall’s $\tau>0.5$ and Spearman’s $\rho>0.65$, indicating a strong alignment across all segments.
@@
Figure 3. Query-level $\boldsymbol{sDCG@K}$ error distribution for single group (left) and paired differences (right) in US market relevance evaluation.
[…]
Query-level $\boldsymbol{sDCG@K}$ error distribution for single group (left) and paired differences (right) in US market relevance evaluation.
@@ 4.2. RQ2: Metrics Sensitivity and MDEs
Before the introduction of LLM labeling, relevance measurement had large MDEs, typically ranging from 1.3% to 1.5%. These large MDEs were primarily the result of the constraints on our sampling designs imposed by the high cost and time consumption of human labeling. The introduction of LLM labeling enabled us to redesign our sampling approach. We increased our sample sizes, moved from simple random sampling (SRS) to stratified sampling, and now use a stratified sampling estimator (Thompson, 2012). Optimal allocation (Neyman, 1934) is used to allocate sample units to strata, which are defined as the intersection of query interest and popularity segment. These changes enabled us to reduce our MDEs to $\leq 0.25\%$.
To attribute the MDE reduction to each source of change, we follow a standard derivation of MDE (Bloom, 1995) as a function of the number of queries in the sample ($n$), metric mean ($\hat{\mu}$), and variance ($\hat{\sigma}^{2}$) assuming standard values for $\alpha=0.05$ and $\beta=0.8$,
@@
Stratified Stratified 2000 0.011 94%
Stratified SRS 5000 0.117 36%
Stratified Stratified 5000 <0.001<0.001 99.6%
@@ 4.3. RQ3: Performance on Non-English Queries
Figure 4. Query-level $\boldsymbol{sDCG@K}$ error distribution for single group (left) and paired differences (right) in France (top) and Germany (bottom) markets relevance evaluation.
[…]
Query-level $\boldsymbol{sDCG@K}$ error distribution for single group (left) and paired differences (right) in France (top) and Germany (bottom) markets relevance evaluation.

### v2

@@ 1. Introduction
Relevance evaluation typically relies on human annotations. However, this approach suffers from high costs, long turnaround time, and limited scalability, which can ultimately compromise the quality of relevance metrics. In recent years, with the advancement of Large Language Models (LLMs), there is a growing interest in automating relevance assessment in the Information Retrieval (IR) community. There have been several works exploring the feasibility of using LLMs to supplement laborious human labeling efforts (faggioli2023perspectives; thomas2024large; upadhyay2024umbrela).
@@ 2.1. Search Relevance Modeling
For LLM-based search relevance modeling, there are two types of architectures commonly used in practice: bi-encoders and cross-encoders. Bi-encoder models encode queries and documents separately into a common dense space and then score their relevance using vector dot-product or cosine similarity. These models are typically trained by minimizing the contrastive loss with in-batch negative sampling (karpukhin2020dense; ma2023fine). While bi-encoder models are efficient for retrieval tasks, they are limited in their ability to capture complex interactions between queries and documents. In contrast, cross-encoder models encode queries and documents jointly, hence can better capture the interaction between queries and documents. Various loss functions are used in the literature for training cross-encoder relevance models, such as pointwise classification loss (nogueira2020document), pairwise or listwise ranking loss (zhuang2023rankt5), and contrastive loss (ma2023fine). In this work, we adopt a cross-encoder architecture and fine-tune the LLMs by minimizing the point-wise multi-class classification loss. The fine-tuned LLMs are then applied to whole-page relevance assessment at Pinterest Search.
@@ 2.2. LLMs for Relevance Judgments
Recent studies have shown the potential of using LLMs for automating relevance assessments in Information Retrieval (IR), as part of the broader “LLM-as-a-judge” paradigm (gu2024survey; zheng2023judging). Initial explorations by faggioli2023perspectives examined the application of LLMs to relevance judgment, providing insights into human-LLM collaboration. Subsequent work by thomas2024large and upadhyay2024umbrela further demonstrated the capability of LLMs to accurately predict searchers’ preferences and generate human-quality relevance labels. Most current research investigates various prompting techniques for LLM-based relevance labeling, including zero-shot (upadhyay2024umbrela; arabzadeh2025benchmarking; upadhyay2024large) and few-shot (macavaney2023one; pires2025expanding) approaches, with a primary focus on evaluating their alignment with human labeling. Later, meng2025query and abbasiantaeb2024can showed that fine-tuning open-source LLMs on human-labeled relevance judgments is crucial for obtaining more reliable relevance predictions, suggesting that fine-tuning much smaller LLMs can yield more effective relevance prediction than few-shot prompting with significantly larger models.
Unlike most existing work that primarily compares different LLMs and prompting techniques (meng2025query; abbasiantaeb2024can) or explores human-LLM collaboration (shankar2024validates), our work presents a real-world application of LLM-based relevance assessment to measure relevance change in online experiments at Pinterest Search. We discuss the practical considerations and validation steps, and highlight the opportunities for improving relevance measurement driven by this approach.
@@ 3.2. Fine-tuned LLMs as Relevance Model
Figure 1. The cross-encoder architecture in the relevance teacher model. Take the encoder language models (e.g., BERT-based models) for illustration.:
Figure 2. Components of LLM-based relevance measurement at Pinterest Search.:
To effectively represent each Pin for relevance prediction, we leverage a comprehensive set of textual features, including Pin titles and descriptions, BLIP image captions (li2022blip), linked page titles and descriptions, user-curated board titles where the Pin has been saved, and highly-engaged queries associated with the Pin. These features together form a robust text representation crucial for accurate relevance assessment.
We experiment with various pre-trained language models, including multilingual BERTbase (devlin2018bert), T5base (2020t5), mDeBERTaV3base (he2021debertav3), XLM-RoBERTalarge (conneau2019unsupervised), and Llama-3-8B (llama3modelcard). The comparative performance for these models and ablation studies on Pin text features can be found in wang2024improving. We then use the fine-tuned model to generate 5-dimensional relevance scores and use the label corresponding to the highest score (argmax) for relevance assessment.
@@ 3.3. Stratified Sampling Design
Stratification plays an important role in sampling-based measurement. First, stratification ensures the sample population is representative of the whole population. In addition, if the strata are chosen such that each stratum is relatively homogeneous, variance reduction can be achieved (miratrix2013adjusting). From a statistical perspective, consider the whole-page relevance score as a random variable $Y$ with mean $\mu$ and variance $\sigma^{2}$. The variance of the sample mean $\bar{Y}$ is given by $V(\bar{Y})=\sigma^{2}/N$, where $N$ is the sample size. If $K$ strata are chosen such that each stratum has a different mean, $V(\bar{Y})$ can be decomposed into within-strata and between-strata variance (see Equation 1), the latter of which can be removed due to stratified sampling (xie2016improving),
@@
To determine the query strata, we evaluated multiple strata choices. While BERTopic (grootendorst2022bertopic) produced the largest variance reduction, we adopted the in-house query-to-interest model based on DistilBERT (sanh2019distilbert), due to ease of integration with existing systems and scalability.
@@ 3.4. Relevance Measurement with LLMs
To measure the relevance impact of an A/B experiment on search ranking, we take a stratified sample of paired search queries from control and treatment experiment groups, ensuring that the sample is representative of overall user usage (thompson2012sampling). The use of paired samples blocks between-query differences, an important source of variation in experiment measurement.
@@
(2) $sDCG@K=\frac{\sum_{k=1}^{K}L_{k}/log_{2}(1+k)}{\sum_{k=1}^{K}5/log_{2}(1+k)},\text{ }L_{k}\in\{1,2,3,4,5\}.$
Lastly, we calculate heterogeneous effects by query popularity and query interest (e.g. beauty, women’s fashion, art, etc), utilizing a Benjamini-Hochberg procedure (benjamini1995controlling) to control the false discovery rate. The LLM-based relevance measurement procedure at Pinterest Search is illustrated in Figure 2.
@@ 4.1. RQ1: Alignment with Human Labels
Following the established best practices in literature (meng2025query; abbasiantaeb2024can) to measure alignment between LLMs and human labels, we compute and report the rank-based correlation Kendall’s $\tau$ and Spearman’s $\rho$ to assess the correlation between the two rankings at query-level sDCG@K metric. To understand the performance on queries with different popularity, we also categorize the queries into 4 popularity segments based on search volume (jain2010organizing): head, torso, tail, and single [^2] Single queries are defined as those that receive fewer than 10 searches.. The results are summarized in Table 1. For all query popularity segments, we achieve Kendall’s $\tau>0.5$ and Spearman’s $\rho>0.65$, indicating a strong alignment across all segments.
@@
Figure 3. Query-level $\boldsymbol{sDCG@K}$ error distribution for single group (left) and paired differences (right) in US market relevance evaluation. :
@@ 4.2. RQ2: Metrics Sensitivity and MDEs
Before the introduction of LLM labeling, relevance measurement had large MDEs, typically ranging from 1.3% to 1.5%. These large MDEs were primarily the result of the constraints on our sampling designs imposed by the high cost and time consumption of human labeling. The introduction of LLM labeling enabled us to redesign our sampling approach. We increased our sample sizes, moved from simple random sampling (SRS) to stratified sampling, and now use a stratified sampling estimator (thompson2012sampling). Optimal allocation (neyman_allocation) is used to allocate sample units to strata, which are defined as the intersection of query interest and popularity segment. These changes enabled us to reduce our MDEs to $\leq 0.25\%$.
To attribute the MDE reduction to each source of change, we follow a standard derivation of MDE (ds_power) as a function of the number of queries in the sample ($n$), metric mean ($\hat{\mu}$), and variance ($\hat{\sigma}^{2}$) assuming standard values for $\alpha=0.05$ and $\beta=0.8$,
@@
Stratified Stratified 2000 0.096 52%
Stratified SRS 5000 0.094 51%
Stratified Stratified 5000 0.061 67%
@@ 4.3. RQ3: Performance on Non-English Queries
Figure 4. Query-level $\boldsymbol{sDCG@K}$ error distribution for single group (left) and paired differences (right) in France (top) and Germany (bottom) markets relevance evaluation.: